# DoubleTake: Geometry Guided Depth Estimation

### v1

@@ Abstract
Estimating depth from a sequence of posed RGB images is a fundamental computer vision task, with applications in augmented reality, path planning etc. Prior work typically makes use of previous frames in a multi view stereo framework, relying on matching textures in a local neighborhood. In contrast, our model leverages historical predictions by giving the latest 3D geometry data as an extra input to our network. This self-generated geometric hint can encode information from areas of the scene not covered by the keyframes and it is more regularized when compared to individual predicted depth maps for previous frames. We introduce a Hint MLP which combines cost volume features with a hint of the prior geometry, rendered as a depth map from the current camera location, together with a measure of the confidence in the prior geometry. We demonstrate that our method, which can run at interactive speeds, achieves state-of-the-art estimates of depth and 3D scene reconstruction in both offline and incremental evaluation scenarios.
@@ 1 Introduction
High quality depth estimations have been shown to be effective for virtual occlusions [67], path planning [14], object avoidance and a variety of augmented reality (AR) effects [14]. For all these applications, we need per-frame depths to be generated at interactive speeds. While the best quality depth maps can be achieved via offline methods e.g. [28, 75], these are not suitable for interactive applications. For interactive use, depths are typically estimated via a multi-view stereo (MVS) approach, where a network takes as input a target frame together with nearby posed and matchable source frames to build a cost volume [31] and estimate a depth map as the output of a neural network [54, 5, 77, 78, 7].
@@
Our contributions are:
@@ 2 Related Work
Depth estimation from posed monocular videos is a long-standing problem in computer vision. Traditional patch-based solutions [17, 56] have been outpaced by learning strategies [74, 26], where the cost volume is generally built by warping features from multiple source frames at different depth hypotheses [10]. The resulting 4D volume is frequently regularized by 3D convolutions [74], which are generally expensive in terms of memory. To tackle this, pyramidal approaches [73, 20] compute multiple cost volumes at different resolutions with a narrow set of hypotheses, and depths computed at lower scales can provide geometric priors to following computation [79, 3]. SimpleRecon [54] shows that accurate depths can be achieved only via 2D convolutions, using cost volumes enriched by additional metadata, while FineRecon [62] further improves the results with a resolution-agnostic 3D training loss and a depth guidance strategy. These multi-view stereo (MVS) approaches have several failure cases: (i) unmatched surfaces, e.g. empty areas in the cost volume where no source frames have a matching view; (ii) depth planes are sparser at greater depths; (iii) greater depths require wider baselines and many more keyframes. Our proposed approach helps overcome these problems by injecting prior information into the network based on previously computed geometry data.
@@
For example, [8, 37, 60] warp previous predictions or features forward and use these as input to the current prediction. Song et al. [60] input previous depth estimates together with a confidence estimate. We also use an estimate of confidence, but our geometry priors are from a global reconstruction, enabling priors from long before in time. Instead of regularizing the cost volume, CER-MVS [43] iteratively updates a disparity field using recurrent units. Recurrent layers [15] and Gaussian Processes [24] have also been used to ingest sequence data. Our work uses sequences but we do not rely on generic layers to encode the prior information. Instead, we directly extract prior knowledge from a 3D reconstruction of the scene to guide the model. An alternative use of sequences is to update predicted depth maps or network weights at test time by optimizing image reconstruction losses e.g. [4, 6, 40, 44, 58, 35, 28] . These methods require the entire sequence to be seen ahead of time, and aren’t suitable for interactive applications.
@@
When available, additional knowledge about the scene might be injected to boost depth estimation methods. In autonomous driving [19, 45], for instance, LiDAR scans are often completed into a dense depth map [41, 42, 70, 9, 70]. Similarly, sparse point clouds from Simultaneous Localization And Mapping (SLAM) algorithms can be used as input to improve depth estimators [71, 72]. However, these priors might be unevenly scattered or unavailable. To tackle these issues, sparsity-agnostic methods [65, 11] and networks capable of both predicting and completing depth maps [22] have been proposed. mvgMVS [49] uses depth data from a depth sensor to directly modulate values in the cost volume. In contrast, our approach uses prior geometry generated from our own estimates as an additional input to an MLP alongside the cost volume. Khan et al. [33] convert per-frame depth estimates to temporally consistent depth via a globally fused point cloud which is input into their final depth network. Their focus is primarily temporal consistency; we compare with [33] and find our approach achieves higher quality depths. Like us, 3DVNet [52] maintain a global reconstruction to improve depth maps. However, their reconstruction and depths are iteratively refined, precluding online use. We also employ priors to improve current depth predictions. Unlike these prior works, we argue that previous predictions of the model itself, if properly managed, are strong hints to improve current estimates for online depth estimation. Furthermore, we also show that our model is robust even in the absence of such information.
@@
Obtaining a 3D reconstruction of a scene is vital for several high-level applications. Estimated depths (e.g. from MVS) can be fused into a TSDF to extract a mesh via Marching Cubes [39]. On the other hand, feedforward volumetric methods [47, 63, 18] directly estimate the volume occupancy (usually encoded as a TSDF), often leveraging expensive 3D CNNs. Implicit representations [48, 76, 36] can generate high quality reconstructions via test-time optimization. These solutions are generally computationally expensive as they typically require a per-scene optimization. Neural Radiance Fields (NeRF) [46] and Gaussian Spatting [32] enable the synthesis of novel views, but their meshes tend to be noisy and additional post-processing steps [50, 21] or a different scene representation [34] might be necessary to generate consistent 3D representations. Extensions to NeRFs utilize sparse structure-from-motion (SfM) point clouds and monocular depth estimation for improved view synthesis [53, 13, 66] or reconstructions [16, 76]. These offline approaches either (a) require the whole video to be seen ahead of time, and/or (b) require compute time to process, making them unsuitable for interactive applications e.g. AR effects and path planning. Conversely, our proposed strategy efficiently estimates depth maps using a 2D MVS model boosted by its own predictions, unlocking accurate reconstructions and depth estimates with a low overhead in computation.
@@ 3 Method
Similar to previous MVS methods e.g. [74, 54], we rely on features matched between the current frame and a few recent source frames to build a feature-based cost volume. Our key contribution is to show how a 3D representation of the scene, constructed from previous depth estimates, can be used as additional input to our network to improve depth predictions for each new frame. In particular, this geometric ‘hint’ complements the likelihood depth estimated via multi-view-stereo on the current and source frames. We incorporate this extra source of information by combining it with the information provided by the cost volume using a “Hint MLP.” An overview of our method is in Fig. 2.
3.1 MVS background
Our method builds on SimpleRecon [54], a state-of-the-art depth estimation method from posed images. Similar to other MVS methods e.g. [10, 31, 74], SimpleRecon constructs a cost volume using multiple viewpoints of the same scene. To compute the cost volume, SimpleRecon starts by extracting feature maps for the current image as well as associated source frames. Extracted source-frame features are then warped to the current camera at each hypothesis depth plane and compared against features for the current frame. This comparison, together with additional metadata, produces a $C{\times}D{\times}H{\times}W$, feature volume with depth likelihood estimates at each location, where $C$ is the channel dimension, $D$ the number of depth planes and $H{\times}W$ the dimensions of the feature maps. A matching MLP is applied to each aggregated feature vector in parallel to obtain the final cost volume with dimensions $D{\times}H{\times}W$. The advantage of this reduction step is that the cost volume can then be processed using 2D convolutions, in contrast to methods that bypass this step and require expensive 3D convolutions. A follow-up monocular prior head regularizes the volume using 2D CNNs and a depth decoder finally produces the output depth map.
3.2 Hint MLP
In addition to using source and target frames, our method allows the input of a rendered depth and a confidence map from a prior geometric representation of the scene. Note that this geometry could be from the distant past, but it could also be from the recent past e.g. as constructed incrementally from previous frames we have recently seen. The rendered depth map is a 2D image of depth values, where each pixel represents the rendered depth from the current camera position to our prior estimate of geometry along that camera ray. The confidence map is a 2D image where each pixel gives an estimate of how certain we are that the prior depth estimate is correct at that pixel, where values of $1.0$ indicate that we have extremely high confidence in the prior depth at this pixel and values of $0.0$ mean a very low confidence in the rendered geometry at this pixel.
Following [54] we create a feature volume and apply a matching MLP to each combined feature vector individually to give a matching score, creating a cost volume with dimensions $D{\times}H{\times}W$. Inspired by this, we use an additional Hint MLP to combine the information from the cost volume with the rendered depth and confidence images. Like the matching MLP, the Hint MLP is applied in parallel at each spatial location and depth plane in the feature volume. Our Hint MLP takes as input a vector with three elements: (i) the matching score from the cost volume, (ii) the geometry hint, defined as the absolute difference between the TSDF rendered depth and the current depth plane and (iii) the confidence value at that pixel. For pixels where there isn’t a rendered depth value, for example because they haven’t been seen before, we set the confidence to 00 and the geometry hint to $-1$. This is done both at training and test time.
See Fig. 3 for an illustration of how the Hint MLP is incorporated into the network architecture. We validate our choice of how to combine the different sources of information in our ablations in Section 4.3.
@@ 3.3 Maintaining persistent 3D geometry
Rendered depth and rendered confidence.
For each new frame, given the corresponding intrinsics and extrinsics, we render a depth map and a confidence map from the TSDF from the point of view of the camera. This is achieved using marching cubes [39] to obtain a mesh, followed by a mesh rendering step. We render the mesh using PyTorch3D [51], which runs on the GPU. This mesh rendering is quick to run, taking only 9.2ms per render on average (see Section 4.2 for more timings). We obtain the confidence map by sampling the voxel confidence channel in the TSDF with coordinates from the backprojected rendered depth; this takes less than 1ms.
Updating the persistent 3D geometry.
Once we have estimated a depth map for a new frame we aggregate it into the current geometry estimate. Our fusion follows InfiniTAM [30, 29]. When updating the TSDF values we also update the confidence measures. For this we use [30, 29]’s quadradtic falloff based on the distance from each voxel to the current camera position, which we clamp to a minimum value of 0.25 to ensure even distant predictions within a maximum assigned distance have some confidence.
ScanNetV2 7Scenes
[…]
DPSNet [27] .1552 .0795 .0299 49.36 93.27 .1966 .1147 .0550 38.81 87.07
MVDepthNet [69] .1648 .0848 .0343 46.71 92.77 .2009 .1161 .0623 38.81 87.70
DELTAS [59] .1497 .0786 .0276 48.64 93.78 .1915 .1140 .0490 36.36 88.13
GPMVS [24] .1494 .0757 .0292 51.04 93.96 .1739 .1003 .0462 42.71 90.32
DeepVideoMVS, fusion [15]* .1186 .0583 .0190 60.20 96.76 .1448 .0828 .0335 47.96 93.79
SimpleRecon [54] .0873 .0430 .0128 74.12 98.05 .1045 .0575 .0156 60.12 97.33
Table 1: Depth evaluation on ScanNetV2 and 7Scenes. Unless stated otherwise, predictions are computed incrementally, without access to future frames or frames from previous scans of the scene. We highlight the best, second-best and third-best methods per metric. Previous scores are from [15, 54]. *[15] was boosted by computing depths using three inference frames instead of two; they also use a custom 90/10 split.
@@ 3.4 Sources of prior geometry
    Alternatively, when we revisit a location we have previously seen, we can load previously generated geometry to use for geometry hints. In this scenario, we assume the current camera position is in the same coordinate system as the loaded geometry. However, we may find that some items have moved since the original geometry was created. Our experiments, with the 3RScan dataset [68], evaluate precisely this scenario.
@@ 3.5 Training data generation
For 50% of the training items we don’t provide a geometry hint, instead feeding $-1$ as the rendered depth and 00 as the confidence; in this scenario, the network must learn to rely only on the matching cost. For the other 50% of the training items the network has access both to the matching cost and to the rendered depth and rendered confidence. We generate a dataset of geometry hints by fusing depths from [54] into a training-time TSDF. 50% of the time that we give a training-time hint it is rendered from the full and complete TSDF, assuming the entire scene has been previously fused, and 50% a partial TSDF fusing only frames up to the current training frame.
Input image Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SR [54] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DVMVS [15] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative depth results on ScanNetV2. All methods are run incrementally, where we run at interactive speeds only with access to previous frames. We compare with [54] and [15], the two closest-performing baselines. Our depth maps are more accurate with better small details (e.g. top) and overall geometry (e.g. bottom).

### v2

@@ Abstract
Estimating depth from a sequence of posed RGB images is a fundamental computer vision task, with applications in augmented reality, path planning etc. Prior work typically makes use of previous frames in a multi view stereo framework, relying on matching textures in a local neighborhood. In contrast, our model leverages historical matches by giving the latest 3D geometry data as an extra input to our network. This self-generated geometric hint can encode information from areas of the scene not covered by the keyframes and it is more regularized when compared with the predicted depth maps for previous frames. We introduce a Hint MLP which combines cost volume features with a hint of the prior geometry, rendered as a depth map from the current camera location, together with a measure of the confidence in the prior geometry. We demonstrate that our method, which can run at 12fps, achieves state-of-the-art estimates of depth and 3D scene reconstruction in both offline and incremental evaluation scenarios.
@@ 1 Introduction
High quality depth estimations have been shown to be effective for virtual occlusions [valentin2018depth], path planning [du2020depthlab], object avoidance and a variety of augmented reality (AR) effects [du2020depthlab]. For all these applications, we need per-frame depths to be generated at interactive speeds. While the best quality depth maps can be achieved via offline methods e.g. [izquierdo2023sfm, yariv2021volume], these are not suitable for interactive applications. For interactive use, depths are typically estimated via a multi-view stereo (MVS) approach, where a network takes as input a target frame together with nearby posed and matchable source frames to build a cost volume [kendall2017end] and estimate a depth map as the output of a neural network [sayed2022simplerecon, chang2018pyramid, Zhang2019GANet, zhang2019domaininvariant, cheng2019learning].
@@
We will release code on acceptance, and our contributions are:
@@ 2 Related Work
Depth estimation from posed monocular videos is a long-standing problem in computer vision. Traditional patch-based solutions [furukawa2015multi, schonberger2016pixelwise] have been outpaced by learning strategies [yao2018mvsnet, huang2018deepmvs], where the cost volume is generally built by warping features from multiple source frames at different depth hypotheses [collins1996space]. The resulting 4D volume is frequently regularized by 3D convolutions [yao2018mvsnet], which are generally expensive in terms of memory. To tackle this, pyramidal approaches [yang2020cost, gu2020cascade] compute multiple cost volumes at different resolutions with a narrow set of hypotheses, and depths computed at lower scales can provide geometric priors to following computation [zhang2023geomvsnet, cai2023riav]. SimpleRecon [sayed2022simplerecon] shows that accurate depths can be achieved only via 2D convolutions, using cost volumes enriched by additional metadata, while FineRecon [Stier_2023_ICCV] further improves the results with a resolution-agnostic 3D training loss and a depth guidance strategy. These multi-view stereo (MVS) approaches have several failure cases: (i) unmatched surfaces, e.g. empty areas in the cost volume where no source frames have a matching view; (ii) depth planes are sparser at greater depths; (iii) greater depths require wider baselines and many more keyframes. Our proposed approach helps overcome these problems by injecting prior information into the network based on previously seen geometry. [GB: based on previously computed geometry data?]
@@
For example, [Stereo-Matching-in-cheng-2023, Raft-stereo-Multilevel-lipson-2021, song2023prior] warp previous predictions or features forward and use these as input to the current prediction. Song et al. [song2023prior] input previous depth estimates together with a confidence estimate. We also use an estimate of confidence, but our geometry priors are from a global reconstruction, enabling priors from long before in time. Instead of regularizing the cost volume, CER-MVS [ma2022multiview] iteratively updates a disparity field using recurrent units. Recurrent layers [duzceker2021deepvideomvs] and Gaussian Processes [hou2019multi] have also been used to ingest sequence data. Our work uses sequences but we do not rely on generic layers to encode the prior information. Instead, we directly extract prior knowledge from a 3D reconstruction of the scene to guide the model. An alternative use of sequences is to update predicted depth maps or network weights at test time by optimizing image reconstruction losses e.g. [casser2018depth, chen2019self, luo2020consistent, mccraith2020monocular, shu2020feature, kuznietsov2021comoda, izquierdo2023sfm] . These methods require the entire sequence to be seen ahead of time, and aren’t suitable for interactive applications.
@@
When available, additional knowledge about the scene might be injected to boost depth estimation methods. In autonomous driving [geiger2012we, Menze2015CVPR], for instance, LiDAR scans are often completed into a dense depth map [ma2018sparse, ma2019self, Depth-Completion-using-Du-2022, choe2021volumetric, Depth-Completion-using-Du-2022]. Similarly, sparse point clouds from Simultaneous Localization And Mapping (SLAM) algorithms can be used as input to improve depth estimators [wong2021unsupervised, Xin2023ISMAR]. However, these priors might be unevenly scattered or unavailable. To tackle these issues, sparsity-agnostic methods [uhrig2017sparsity, conti2023sparsity] and networks capable of both predicting and completing depth maps [guizilini2021sparse] have been proposed. mvgMVS [poggi2022multi] uses depth data from a depth sensor to directly modulate values in the cost volume. In contrast, our approach uses prior geometry generated from our own estimates as an additional input to an MLP alongside the cost volume. Khan et al. [khan2023temporally] convert per-frame depth estimates to temporally consistent depth via a globally fused point cloud which is input into their final depth network. Their focus is primarily temporal consistency; we compare with [khan2023temporally] and find our approach achieves higher quality depths. Like us, 3DVNet [rich20213dvnet] maintain a global reconstruction to improve depth maps. However, their reconstruction and depths are iteratively refined, precluding online use. We also employ priors to improve current depth predictions. Unlike these prior works, we argue that previous predictions of the model itself, if properly managed, are strong hints to improve current estimates for online depth estimation. Furthermore, we also show that our model is robust even in the absence of such information.
@@
Obtaining a 3D reconstruction of a scene is vital for several high-level applications. Estimated depths (e.g. from MVS) can be fused into a TSDF to extract a mesh via Marching Cubes [lorensen1998marching]. On the other hand, feedforward volumetric methods [murez2020atlas, sun2021neuralrecon, gao2023visfusion] directly estimate the volume occupancy (usually encoded as a TSDF), often leveraging expensive 3D CNNs. Implicit representations [peng2020convolutional, yu2022monosdf, li2023neuralangelo] can generate high quality reconstructions via test-time optimization. These solutions are generally computationally expensive as they typically require a per-scene optimization. Neural Radiance Fields (NeRF) [mildenhall2020nerf] and Gaussian Spatting [kerbl20233d] enable the synthesis of novel views, but their meshes tend to be noisy and additional post-processing steps [rakotosaona2023nerfmeshing, guedon2023sugar] or a different scene representation [Kulhanek_2023_ICCV] might be necessary to generate consistent 3D representations. Extensions to NeRFs utilize sparse structure-from-motion (SfM) point clouds and monocular depth estimation for improved view synthesis [roessle2022dense, kangle2021dsnerf, scade-nerfs-from-Uy-2023] or reconstructions [fu2022geo, yu2022monosdf]. These offline approaches either (a) require the whole video to be seen ahead of time, and/or (b) require compute time to process, making them unsuitable for interactive applications e.g. AR effects and path planning. Conversely, our proposed strategy efficiently estimates depth maps using a 2D MVS model boosted by its own predictions, unlocking accurate reconstructions and depth estimates with a low overhead in computation.
@@ 3 Method
Similar to previous MVS methods e.g. [yao2018mvsnet, sayed2022simplerecon], we rely on features matched between the current frame and a few recent source frames to build a feature-based cost volume. Our key contribution is to show how a 3D representation of the scene, constructed from previous depth estimates, can be used as additional input to our network to improve depth predictions for each new frame. In particular, this geometric ‘hint’ complements the likelihood depth estimated via multi-view-stereo on the current and source frames. We incorporate this extra source of information by combining it with the information provided by the cost volume using a “Hint MLP.” An overview of our method can be seen in Fig. 2.
3.1 MVS background
Our method builds on SimpleRecon [sayed2022simplerecon], a state-of-the-art depth estimation method from posed images. Similar to other MVS methods e.g. [collins1996space, kendall2017end, yao2018mvsnet], SimpleRecon constructs a cost volume using multiple viewpoints of the same scene. To compute the cost volume, SimpleRecon starts by extracting feature maps for the current image as well as associated source frames. Extracted source-frame features are then warped to the current camera at each hypothesis depth plane and compared against features for the current frame. This comparison, together with additional metadata, produces a $C{\times}D{\times}H{\times}W$, feature volume with depth likelihood estimates at each location, where $C$ is the channel dimension, $D$ the number of depth planes and $H{\times}W$ the dimensions of the feature maps. A matching MLP is applied to each aggregated feature vector in parallel to obtain the final cost volume with dimensions $D{\times}H{\times}W$. The advantage of this reduction step is that the cost volume can then be processed using 2D convolutions, in contrast to methods that bypass this step and require expensive 3D convolutions. A follow-up monocular prior head regularizes the volume using 2D CNNs and a depth decoder finally produces the output depth map.
3.2 Hint MLP
In addition to using source and target frames, our method allows the input of a rendered depth and a confidence map from a prior geometric representation of the scene. Note that this geometry could be from the distant past, but it could also be from the recent past e.g. as constructed incrementally from previous frames we have recently seen. The rendered depth map is a 2D image of depth values, where each pixel represents the rendered depth from the current camera position to a prior estimate of geometry along that camera ray. The confidence map is a 2D image where each pixel gives an estimate of how certain we are that the prior depth estimate is correct at that pixel, where values of $1.0$ indicate that we have extremely high confidence in the prior depth at this pixel and values of $0.0$ mean a very low confidence in the rendered geometry at this pixel.
Following [sayed2022simplerecon] we create a feature volume and apply a matching MLP to each combined feature vector individually to give a matching score, creating a cost volume with dimensions $D{\times}H{\times}W$. Inspired by this, we use an additional hint MLP to combine the information from the cost volume with the rendered depth and confidence images. Like the matching MLP, the Hint MLP is applied in parallel at each spatial location and depth plane in the feature volume. Our Hint MLP takes as input a vector with three elements: (i) the matching score from the cost volume, (ii) the geometry hint, defined as the absolute difference between the TSDF rendered depth and the current depth plane and (iii) the confidence value at that pixel. For pixels where there isn’t a rendered depth value, for example because they haven’t been seen before, we set the confidence to 00 and the geometry hint to $-1$. This is done both at training and test time.
@@ 3.2 Hint MLP
See Fig. 3 for an illustration of how the Hint MLP is incorporated into the network architecture. We validate our choice of how to combine the different sources of information in an ablation study in our ablations in Section 4.3.
@@ 3.3 Maintaining persistent 3D geometry
Rendered depth and rendered confidence
For each new frame, given the corresponding intrinsics and extrinsics, we render a depth map and a confidence map from the TSDF from the point of view of the camera. This is achieved using marching cubes [lorensen1998marching] to obtain a mesh, followed by a mesh rendering step. We render the mesh using PyTorch3D [ravi2020pytorch3d], which runs on the GPU. This mesh rendering is quick to run, taking only 6.9ms per render on average (see Section 4.2 for more timings). We obtain the confidence map by sampling the voxel confidence channel in the TSDF with coordinates from the backprojected rendered depth, or alternatively, encode the confidence as vertex colors in the mesh. Either takes less than 1ms.
Updating the persistent 3D geometry
Once we have estimated a depth map for a new frame we aggregate it into the current geometry estimate. Our fusion follows InfiniTAM [InfiniTAM_ECCV_2016, InfiniTAM_ISMAR_2015]. When updating the TSDF values we also update the confidence measures. For this we use [InfiniTAM_ECCV_2016, InfiniTAM_ISMAR_2015]’s quadradtic falloff based on the distance from each voxel to the current camera position, which we clamp to a minimum value of 0.25 to ensure even distant predictions within a maximum assigned distance have some confidence assigned.
@@ 3.4 Sources of prior geometry
    Alternatively, when we revisit a location we have previously seen, we can load previously generated geometry to use for geometry hints. In this scenario, we assume the current camera position is in the same coordinate system as the loaded geometry. However, we may find that some items have moved since the original geometry was created. Our experiments, with the 3RScan dataset [Wald2019RIO], evaluate precisely this scenario.
@@ 3.5 Training data generation
For 50% of the training items we don’t provide a geometry hint, instead feeding $-1$ as the rendered depth and 00 as the confidence; in this scenario, the network must learn to rely only on the matching cost. For the other 50% of the training items the network has access both to the matching cost and to the rendered depth and rendered confidence. We generate a dataset of geometry hints by fusing depths from [sayed2022simplerecon] into a training-time TSDF. 50% of the time that we give a training-time hint it is rendered from the full and complete TSDF, assuming the entire scene has been previously fused, and 50% a partial TSDF fusing only frames up to the current training frame.
Figure 4: We introduce a more accurate mesh evaluation. (a) shows the ground truth mesh, which contains many holes. (b) shows an example predicted mesh, here from [Stier_2023_ICCV]. This is punished for being too complete, as [bozic2021transformerfusion]’s visibility mask (c) extends beyond the ground truth, giving high Acc error on their prediction (d). Our new masking (e) is tighter to the ground truth mesh, giving a more meaningful error (f).
3.6 Implementation details
We train our model (and all our ablations) with a batch size of 16 on each of two Nvidia A100 GPUs. We use a learning rate of $41e{-}$ 4 that we drop by a factor of 10 at $70000$ and $80000$ steps. We train with an input dimension of $512\times 384$, and an output dimension of $256\times 192$. During training, we color augment RGB images. Our network follows that of [sayed2022simplerecon], with the addition of our ‘hint MLP’ as described above. This network uses an EfficientNetV2 S [tan2021efficientnetv2] encoder for the monocular image prior and the first two blocks of a ResNet18 [he2016deep] encoder for generating matching features for the cost volume. The decoder follows UNet++ [zhou2018unet++]. Full architecture details are given in the supplementary material. Our training losses directly follow [sayed2022simplerecon]. We select source frames for building the cost volume using the strategy and hyperparameters from [duzceker2021deepvideomvs]. For evaluation of online methods, we use online source views for each keyframe’s cost volume, so source frames come from the past in a sequence. For evaluation of offline methods, source frames may come from anywhere in a sequence. For both ‘ours’, SimpleRecon [sayed2022simplerecon] and DeepVideoMVS [duzceker2021deepvideomvs] we use the same TSDF fusion method. We fuse depth maps to a maximum depth of $3$m into a TSDF volume of $2$cm resolution. For the FineRecon [Stier_2023_ICCV] baseline we use a $1$cm voxel resolution.
4 Evaluation
We evaluate with three challenging 3D datasets, all acquired with a handheld RGBD sensor. We train and evaluate on ScanNetV2 [dai2017scannet], which comprises 1,201 training scans, 312 validation scans, and 100 testing scans of indoor scenes. We additionally evaluate our ScanNetV2 models on 7-Scenes [shotton2013scene] without fine-tuning, following e.g. [sayed2022simplerecon] and with the test split from [duzceker2021deepvideomvs]. We also evaluate on 3RScan [Wald2019RIO]. This dataset captures the same environment in multiple separate scans, between which objects’ positions have changed. This tests our ability to use scans captured in the past as ‘hints’ for instantaneous depth estimates.
ScanNetV2 7Scenes